THE ELECTRONIC SPECTRUM OF THE NEPTUNYL ION, NpO_{2}^{2+}

\underline{SPIRIDOULA MATSIKA}, RUSSELL M. PITZER , Department of Chemistry, The Ohio State University, 100 W.18th Avenue, Columbus, OH 43210.

The electronic spectrum of the neptunyl ion, NpO_{2}^{2+}, has been calculated using relativistic spin-orbit configurationinteraction calculations. Neptunium in NpO_{2}^{2+} is in oxidation state VI and has an f^{1} configuration. In the lower energies lie the f → f transitions while at energies above 20000 cm^{−1} charge transfer transitions occur. The ground state is a ^{2}Δ state with Ω = 3/2. Equilibrium distances and vibrational frequencies for the symmetric stretch in the ground state are obtained and compared with experimental work.